OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

two paralell lines have same slope..so slope of given line=slope of line whose eq is to be found=-3 so now use this eq. (y-y1)=m(x-x1) here m=slope and (x1,y1)=(4,-2) plug in values...and get result..

OpenStudy (anonymous):

sorrey..slope=+3 not -3..!!!i meesed it up..!!!

OpenStudy (anonymous):

it will be option D..!

OpenStudy (anonymous):

OMG thank you ! :) <3

OpenStudy (yanasidlinskiy):

By the way... \(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)